OpenStudy (anonymous):

subtact -4/19-{-5/19}

OpenStudy (ghazi):

\[\frac{ -4 }{ 19 }-\frac{ -5 }{ 19 }=\frac{ -4-(-5) }{ 19 }=\frac{ -4+5 }{ 19}=\frac{ 1 }{ 19 }\]